OpenStudy (anonymous):

How to do this? Its Slope-Intercept Form. Forgot how to do: What is the equivalent standard form equation for this line? 3y = 6x + 9

OpenStudy (amistre64):

turn it into: Ax +By = C

OpenStudy (anonymous):

How do you do that again? havent don it in awhile and totally forgot how

OpenStudy (amistre64):

you forgot how to subtract 3y from both sides?

OpenStudy (anonymous):

Ohh, no i can do that. I wasnt sure how to change it into ax + by = c well then it would be = 3x + 9 - 3y... right?

OpenStudy (anonymous):

Am i on the right track?..

OpenStudy (anonymous):

Google is saying slope intercept form is y=mx + b? So it would be.. y = 2x + 9?

OpenStudy (anonymous):

Ax + By = C is standard form

OpenStudy (anonymous):

Yes, i know thats the slope interecpt form. I would think that wuld be the anwser but isnt one of the choices :/

OpenStudy (anonymous):

2x - 3y = 9, is that a choice?

OpenStudy (anonymous):

No :/ hate geo

hero (hero):

3y = 6x + 9 Divide both sides by three: y = 2x + 3 Subtract y from both sides/Subtract 3 from both sides: 2x - y = -3

hero (hero):

@LilVeroDez13

OpenStudy (anonymous):

@Hero thanks!

OpenStudy (amistre64):

internet went down as i was typing this: 3y = 6x + 9 if we subtract 3y from both sides we get 6x - 3y + 9 = 0 this is considered standard form in some textbooks, other would have the constant on the other side of the equal sign, so to go further if need be, lets subtract 9 from both sides 6x - 3y = -9 this is also considered standard form, depending on the author of the material you use the divide by 3 is a nice touch tho

OpenStudy (anonymous):

Oh, thanks! thought u jus left.

hero (hero):

I always understood it to be the standard form of the equation of a line converted from the slope-intercept form. Furthermore, I was taught to reduce to simplest form first (if possible) before doing anything else. So that's why I divided both sides by three initially. Either way is correct, but I get the feeling mine is more correct because of the slope-intercept form.

hero (hero):

2x - y = -3 is also standard form wouldn't you agree @amistre64?

OpenStudy (amistre64):

yes, i would agree.

OpenStudy (amistre64):

that would most likely be the simplest form using integer coefficients

OpenStudy (anonymous):

thanks both of you guys :)